# Stable generative modeling using diffusion maps

### v1

@@ 5. Conditional sampling
In other words, we replace the current $X_{n}$ by $\hat{X}_{n}=(Y_{n},z^{\ast})$. Next we apply the split-step scheme (3.2) to $\hat{X}_{n}$, that is,
$X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2\epsilon}K(\hat{X}_{n})^{1/2}\Xi_{n},$
and
$X_{n+1}$ $=m_{\epsilon}(X_{n+1/2})=\mathcal{X}p_{\epsilon}(X_{n+1/2})$
where $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})$ and the definition of the probability vectors $p_{\epsilon}(x)$ follows from (8). The split-step scheme (3.2) generalises along the same lines.
@@ 6. Numerical experiments
In this section, we illustrate our method through three numerical examples encompassing different ranges and focal points. In the first two examples,we generate samples using synthetic datasets with increasing dimensions. Our emphasis is on exploring the impact of employing the variable bandwidth kernel. In the third example, we showcase the proposed conditional generative modeling in Section 5, applied to a stochastic subgridscale parametrization problem.
@@ 6.1. One-dimensional manifold
To illustrate how well the proposed methods generate statistically reliable samples we consider first the case of $M$ samples $x\in{\mathbb{R}}^{2}$. In particular, we consider samples with a polar representation with radius $r=1+\sigma_{r}\xi_{r}$ and angle $\theta=\pi/4+\sigma_{\theta}\xi_{\theta}$ with $\sigma_{r}=0.06$ and $\sigma_{\theta}=0.6$ and $\xi_{r,\theta}\sim{\rm{N}}(0,1)$. We used $M=2,000$ samples to learn the transition kernel and then generated $10,000$ new samples with an initial condition at the tail with the data point corresponding to the smallest angle.
@@
We begin by investigating the effect of the two noise models proposed, namely a constant diffusion as in (13) with constant bandwidth $K(x)=I$ and the case when the diffusion reproduces the sample covariance $C$ as in (16). We employ a Langevin sampler with the splitting scheme (3.2) with $\epsilon=0.009$ and (3.2), respectively. Figure 1 shows that choosing the sample covariance as the noise model is clearly advantageous. Whereas both noise models generate samples that reproduce the angular distribution the noise model using a constant diffusion is overdiffusive in the radial direction. In contrast the noise model using the sample covariance nicely reproduces the radial distribution.
@@
We now investigate the effect of a variable bandwidth $K(x)$. We employ the noise model (3.2) with the sample covariance but use $K(x)$ to determine the diffusion map (cf. (7)). The Langevin sampler (3.2) is again initialized with the coordinates of the data point corresponding to the smallest angle in the data-sparse tail. Figure 2a shows the samples projected onto the convex hull of the data, i.e. outputs of step (19b), when a uniform bandwidth $K(x)=I$ with $\epsilon=0.009$ is employed. Although the mean behaviour is well reproduced, it is seen that the generative model fails near the data-sparse tails for large and small angles. Here the value of $\epsilon$ is too small to generate significant diffusion and the samples are aligned on the (linear) convex hull of the widely separated data samples. To mitigate against this behaviour we employ a variable bandwidth $\rho(x)=\pi^{\beta}$ with $\beta=-1/5$ and a kernel density estimate $\pi(x)$. Figure 2b shows how the variable bandwidth kernel is able to better reproduce the sampling in the data-sparse tail regions. Figure 3 shows the empirical histograms for the radius and the angle variables of the noisy samples corresponding to Figure 2 (i.e. outputs of step (19a)). Whereas both, the uniform and the variable bandwidth kernels, reproduce the radial distribution very well, the uniform bandwidth fails to reproduce the angular distribution in the tails where the diffusion is not sufficiently strong to let the sampler escape the convex hull of the data.
@@
We have seen that a constant uniform bandwidth generates samples which are concentrated in the bulk of the data and which are overly diffusive in the radial direction (cf. Figure 1c). One may wonder if employing a smaller virtual time step $\Delta\tau<\epsilon$ in the Langevin sampler (11) will allow a Langevin sampler with constant bandwidth to generate more faithful samples. Figure 4 shows that choosing a smaller time step $\Delta\tau$ in (15), here with $\Delta\tau=\epsilon/4$ is indeed able to reproduce the radial distribution. However, if the Langevin sampler is initialised with a data point in the center of the data samples, it is not able to diffuse into the tail of the distribution distribution, leading to an under-diffusive empirical histogram for the angles.
@@
Figure 2. Effect of a variable bandwidth $K(x)=\rho(x)I$ in data-sparse regions. For the generative model the Langevin sampler (3.2) is used and we set $\epsilon=0.009$. Results are shown for the output of step (19b). Left: Original (blue) and generated data for a constant bandwidth $K(x)=I$ (red). Right: Original (blue) and generated data for a variable bandwidth $K(x)=\rho(x)I$ with $\rho(x)=\pi(x)^{\beta}$ with $\beta=-1/5$ (magenta).
Figure 3. Effect of a variable bandwidth $K(x)=\rho(x)I$ on the angular and radial distributions (left and right, respectively). Shown are the original data (blue), generated data for a constant bandwidth $K(x)=I$ (red) and for a variable bandwidth $K(x)=\rho(x)I$ with $\rho(x)=\pi(x)^{\beta}$ with $\beta=-1/5$. The data were generated using a constant covariance noise model in (3.2) and $\epsilon=0.009$.
Figure 4. Effect of a variable time step $\Delta\tau$ in the Langevin sampler (3.2) with constant diffusion $K=1$. Results are shown for the original data, and for $\Delta\tau=\epsilon$ and $\Delta\tau=\epsilon/4$. Throughout a constant bandwidth is used. Left: Empirical histogram of the angular variable $\theta$. Middle: Empirical histograms of the radial variable $r$. Right: Original (blue) and generated data in the $(x_{1},x_{2})$-plane with $\Delta\tau=1$ (red) and with $\Delta\tau=\epsilon/4$ (green).
@@ 6.2. Multi-dimensional manifolds
In this numerical example, we show our proposed method on hyper semi-spheres of dimension $d=\{3,4,9\}$, using both a fixed bandwidth kernel and a variable bandwidth kernel. Data are generated by firstly sampling $z^{(i)}=(z_{1}^{(i)},\cdots,z_{d}^{(i)})$ from a $d$-dimensional standard normal distribution, and then setting $y^{(i)}=(z^{(i)}_{1},\cdots,\alpha z^{(i)}_{d})$, with $\alpha=5$ to promote non-uniformity. Finally, the samples $x^{(i)}$ are obtained by normalizing $y^{(i)}$ to achieve the unit length, i.e., $y^{(i)}/\left\|y^{(i)}\right\|$, and perturbing $y^{(i)}/\left\|y^{(i)}\right\|$ in the radial direction with $U(0,0.01)$ noise. An instance of the target samples of three dimensions can be seen in Figure 5. Given a bandwidth $\epsilon$ and a bandwidth function $\rho(x)$, we implement the proposed scheme (3.2). For the fixed bandwidth kernel we set $\rho(x)=1$. For the variable bandwidth kernel (22), we set $\rho(x)=\pi(x)^{\beta}$, with $\beta<0$, and $\pi(x)$ is approximated using a kernel density estimator. We use $M=1,000$ training samples to learn the transition kernel and run a Langevin sampler to generate $50,000$ samples, with the initial data point being $(1,0,\cdots,0)\in\mathbb{R}^{d}$. To obtain a better mixing of the Langevin sampler, we take one every $20$ samples in the the last $2,0000$ generated samples of the chain, resulting in a total of $1,000$ samples. To evaluate the quality of the generated samples, we compute the regularized optimal transport (OT) distance between the generated samples $\{x_{\mathrm{gen}}^{(i)}\}_{i=1}^{M_{\mathrm{g}}}$ and the original reference samples $\{x_{\mathrm{ref}}^{(j)}\}_{j=1}^{M_{\mathrm{r}}}$. The regularized OT distance with entropy penalty $1/\lambda$ is defined as
@@
is the information entropy. The entries of $C\in\mathbb{R}^{M_{g}\times M_{r}}$ are set to be the pairwise Euclidean distances between $\{x_{\mathrm{gen}}^{(i)}\}_{i=1}^{M_{\mathrm{g}}}$ and $\{x_{\mathrm{ref}}^{(j)}\}_{j=1}^{M_{\mathrm{r}}}$, and each sample is assigned equal weight marginally. We compute this distance using the Sinkhorn–Knopp algorithm [9, 26]. The number of reference samples is chosen to be $M_{\mathrm{r}}=50,000$, and the entropic regularization penalty is set to be $1001/\lambda=$ 100. We consider the OT distance as a diagnostic to quantify the statistical accuracy of the sampling scheme.
@@
We then optimize over the parameters $\epsilon$ for a fixed $\lambda$ using grid search. To be more precise, for the Langevin sampler with a fixed bandwidth kernel, we vary $\epsilon$, and compute the OT distance of the generated samples. The best performed $\epsilon$ is chosen to be the one that corresponds to the smallest OT distance, and we call it $\epsilon^{*}$. For the variable bandwidth kernel, we fix $\epsilon=\epsilon^{*}$. In order to disentangle the effect of varying $\epsilon$ and varying $\beta$ we use the normalized variable bandwidth as described in (22). $\beta$ is set to be $-0.01\times 2^{n}$ with $n=\{0,\cdots,8\}$ for $d=\{3,4,9\}$. The optimal bandwidth $\epsilon^{*}$ is reported in Table 1 and the comparisons between the fixed bandwidth kernel and the variable bandwidth kernel are presented in Figure 6. We observe that by keeping $\epsilon$ fixed, the OT distance becomes smaller for a wide range of $\beta$.
@@
We then examine the quality of generated samples at the optimal $\epsilon$ and $\beta$ along the last coordinate (the nonuniform direction). Similar to the previous study, we compute the one dimensional OT distance of the marginal distribution (see Figure 5(b)) and show the histograms and the cumulative density function (CDF) of the samples generated using the fixed bandwidth kernel and using the variable bandwidth kernel in Figure 7. The benefit of using the variable bandwidth kernel becomes more prominent when focusing on the marginal samples. In the case where we sample a 4-dimensional hyper-semisphere with non-uniformity along the last coordinate, we see in Figure 7 that the empirical CDF of the samples generated with the variable bandwidth kernel closely aligns with the reference (constructed using samples from the target distribution) for the most part. In contrast, the samples generated using the fixed variable bandwidth kernel noticeably diverge from the reference. This aligns with what we observed in Figure 2 and Figure 3 — the data generated using the variable bandwidth kernel better resemble the original data. In the cases where the samples are drawn from a 9-dimensional hyper-semisphere, while both methods struggle to generate samples that mirror those from the target distribution, primarily due to the inherent limitations of kernel methods in high-dimensional scenarios, employing a variable bandwidth kernel produces samples that exhibit a closer resemblance to those from the target distribution.
@@
Figure 5. Nonuniform samples $x^{(i)}$ on a 3-dimensional semisphere. The non-uniformity is along the last coordinate $x_{3}$.
@@
(a) (a)
(b) (b)
Figure 6. The dash lines are the OT distance of the samples using a fixed bandwidth kernel and the solid lines are the OT distance of the samples using a variable bandwidth kernel. (a) Comparison between the OT distance of samples generated using a fixed bandwidth kernel and using a variable bandwidth kernel. (b) Comparison between the one-dimensional marginal OT distance of samples generated using a fixed bandwidth kernel and using a variable bandwidth kernel, along the last coordinate (non-uniform direction). Here the Langevin sampler is initialized at $(1,0,\cdots,0)$ for all cases.
@@
6.3. Stochastic subgridscale parametrization
The conditional sampling algorithm described in Section 5 can be used to perform stochastic subgridscale parametrization, a central problem encountered in, for example, the climate sciences. The problem of subgridscale parametrization, or more generally of model closure is the following: given a potentially stiff dynamical system
@@ 6.3. Stochastic subgridscale parametrization
where $\varepsilon<1$ denotes the time scale separation between slow resolved variables of interest $z\in{\mathbb{R}}^{d_{s}}$ and unresolved fast degrees $y\in{\mathbb{R}}^{d_{f}}$. Note the notational difference between the time scale separation parameter $\varepsilon$ and the bandwidth parameter $\epsilon$ used to define the diffusion map. For $\varepsilon\ll 1$ the system is stiff and to ensure numerical stability a small time step $\Delta t<\varepsilon$ is needed. This, together with the potential high-dimensionality of the fast subspace $d_{f}\gg d_{s}$, constitutes a computational barrier for simulating the dynamics on the slow time scale of interest. Hence one is interested in obtaining an effective evolution equation for the slow resolved variables $z$ only which captures the essential effect of the unresolved variables $y$. Hence, we seek to determine the effective reduced dynamics
(27) $\dot{z}=F_{z}(z)+\psi(z),$
where $F_{z}(z)$ denotes some a priori known deterministic drift, possibly based on physical reasoning, and $\psi(z)$ denotes the closure term which may be deterministic or stochastic, and which parametrizes the unknown unresolved fast processes. Deterministic machine learning methods have previously been used to learn the average effect of the unresolved variables, i.e. the average of $g(z,y;\varepsilon)$ over the (conditional) invariant measure of the fast process [18, 19]. Deterministic maps, however, are not able to capture the resolved dynamics with sufficient statistical accuracy, and it is by now well established that the effective equation is of a stochastic nature [34, 17, 23, 16]. In the climate sciences this idea goes back to the seminal work by the 2022 Nobel Prize recipient Klaus Hasselmann who treated the slow ocean as a stochastic dynamical system as a result of it experiencing the integrated effect of fast moving weather systems. In the case of infinite time scale separation there are explicit expressions for the effective drift and diffusion term of the effective slow dynamics. However, these terms include integrals over the auto-correlation functions and are numerically very hard to estimate. Moreover, for the realistic case of moderate time-scale separation Edgeworth corrections need to be included [59, 60], the estimation of which requires even longer time-series. Instead we propose to learn the closure term $\psi(z)$ and generate realisations $\psi(z)$ on the fly employing the conditional sampling algorithm described above. We consider the situation in which scientists have a good understanding of the resolved dynamics and know the slow vector field $F_{z}(z)$. Given data of the resolved variables $\{z_{n}\}_{n=1}^{N}$ sampled at equidistant times $\Delta t$, the closure term $\psi(z)$ capturing the effect of the unresolved dynamics (26b) can then be estimated as
(28) $\psi(z_{n})=z_{n+1}-z_{n}-F_{z}(z_{n})\,\Delta t,$
@@
$z^{\ast}$ $=z_{n}$
$x_{1,j}$ $=z^{\ast}$
$x_{j+1/2}$ $=x_{j}+\sqrt{2\epsilon}\,\Xi_{j}$
$x_{j+1}$ $=\mathcal{Z}p_{\epsilon}(x_{j+1/2}),$
@@
We look at the case of effective additive noise with $h(z)=1$ which does not require conditioning on the slow variable $z$, as well as at the case of multiplicative noise with $h(z)=z$. We used MATLAB’s built-in ode45 routine [32] to generate the time series with a time-scale separation parameter of $\varepsilon=0.01$. The time series is subsequently sub-sampled with $\Delta t=0.1$. Figure 8 shows a comparison between the full multi-scale system (30) - (6.3) with additive noise $h(z)=1$ and the stochastic surrogate model (29). The slow $z$-dynamics exhibits stochastic bimodal dynamics. Figure 9 shows a comparison for the multiplicative case $h(z)=z$ which yields unimodal slow dynamics. It is clearly seen that the surrogate model (29) obtained by the generative conditional sampler generates statistically reliable dynamics.
Figure 8. Results for the stochastic subgrid-scale parametrization for the multi-scale system (30) - (6.3) with $\varepsilon=0.01$ and with additive noise $h(z)=1$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgridscale parametrization scheme using our generative sampler trained with $M=120,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
Figure 9. Results for the stochastic subgrid-scale parametrization for the multi-scale system (30) - (6.3) with $\varepsilon=0.01$ and with multiplicative noise $h(z)=z$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgridscale parametrization scheme using our generative sampler, trained with $M=20,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
7. Conclusions
We introduced a diffusion map based Langevin scheme for generative modeling. Our method combines diffusion maps with discrete-time Langevin-type sampling, yielding a nonparametric generative model that requires minimal tuning. We showed numerically that employing a variable bandwidth kernel, in contrast to a fixed bandwidth kernel, results in generated samples with enhanced accuracy. The performance of the conditional generative model was showcased through its application to a stochastic subgrid-scale parametrization problem. Future research will delve into the theoretical foundations of the proposed scheme, including its convergence rate and scalability.

### v2

@@ 5. Bayesian inference and conditional sampling
In other words, we replace the current $X_{n}$ by $\hat{X}_{n}=(Y_{n},z^{\ast})$. Next we apply the split-step scheme (20) to $\hat{X}_{n}$, i.e.,
$X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2K(\hat{X}_{n})}\Xi_{n},$
$X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2}),$
where we have again $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})$, and the definition of the probability vectors $p_{\epsilon}(x)$ follows from (10). The split-step scheme (21) generalises along the same lines.
@@ 6. Numerical experiments
In this section, we illustrate our method through three numerical examples encompassing different ranges and focal points. In the first two examples, we generate samples using synthetic datasets with increasing dimensions. Our emphasis is on exploring the impact of different diffusions and of employing a variable bandwidth kernel. In the third example, we showcase the proposed conditional generative modeling in Section 5, applied to a stochastic subgrid-scale parametrization problem.
@@ 6.1. One-dimensional manifold
To illustrate how well the proposed methods generate statistically reliable samples we consider first the case of $M$ samples $x\in{\mathbb{R}}^{2}$. In particular, we consider samples with a polar representation with radius $r=1+\sigma_{r}\xi_{r}$ and angle $\theta=\pi/4+\sigma_{\theta}\xi_{\theta}$ with $\sigma_{r}=0.06$ and $\sigma_{\theta}=0.6$ and $\xi_{r,\theta}\sim{\rm{N}}(0,1)$. We used $M=2,000$ samples to learn the transition kernel and then generated $10,000$ new samples with an initial condition from the data-sparse tail of the distribution, chosen to be the data point corresponding to the smallest angle.
@@
We begin by investigating the effect of the two noise models proposed, namely a constant diffusion as in (15) with constant bandwidth $K(x)=I$ and the case when the diffusion reproduces the sample covariance $C$ as in (18). We employ a Langevin sampler with the splitting scheme (20) with $\epsilon=0.009$ and (21), respectively. Figure 1 shows that choosing the sample covariance as the noise model is clearly advantageous. Whereas both noise models generate samples that reproduce the angular distribution the noise model using a constant diffusion is overdiffusive in the radial direction. In contrast the noise model using the sample covariance nicely reproduces the radial distribution.
@@
We now investigate the effect of a variable bandwidth $K(x)$. We employ the noise model (21) with the sample covariance but use $K(x)$ to determine the diffusion map (cf. (9)). The Langevin sampler (21) is again initialized with the coordinates of the data point corresponding to the smallest angle in the data-sparse tail. Figure 2 (left) shows the samples projected onto the convex hull of the data, i.e. outputs of step (21b), when a uniform bandwidth $K(x)=I$ with $\epsilon=0.009$ is employed. Although the mean behaviour is well reproduced, it is seen that the generative model fails near the data-sparse tails for large and small angles. Here the value of $\epsilon$ is too small to generate significant diffusion and the samples are aligned on the (linear) convex hull of the widely separated data samples. To mitigate against this behaviour we employ a variable bandwidth $\rho(x)=\pi^{\beta}$ with $\beta=-1/5$ and a kernel density estimate $\pi(x)$. Figure 2 (right) shows how the variable bandwidth kernel is able to better reproduce the sampling in the data-sparse tail regions. Figure 3 shows the empirical histograms for the radius and the angle variables of the noisy samples corresponding to Figure 2 (i.e. outputs of step (21a)). Whereas both, the uniform and the variable bandwidth kernels, reproduce the radial distribution very well, the uniform bandwidth fails to reproduce the angular distribution in the tails where the diffusion is not sufficiently strong to let the sampler escape the convex hull of the data.
@@
We have seen that a constant uniform bandwidth generates samples which are concentrated in the bulk of the data and which are overly diffusive in the radial direction (cf. Figure 1 (right)). One may wonder if employing a smaller virtual time step $\Delta\tau<\epsilon$ in the Langevin sampler (13) will allow a Langevin sampler with constant bandwidth to generate more faithful samples. Figure 4 shows that choosing a smaller time step $\Delta\tau$ in (17), here with $\Delta\tau=\epsilon/4$ is indeed able to reproduce the radial distribution. However, if the Langevin sampler is initialised with a data point in the center of the data samples, it is not able to diffuse into the tail of the distribution distribution, leading to an under-diffusive empirical histogram for the angles.
@@
Figure 2. Effect of a variable bandwidth $K(x)=\rho(x)I$ in data-sparse regions. For the generative model the Langevin sampler (21) is used and we set $\epsilon=0.009$. Results are shown for the output of step (21b). Left: Original (blue) and generated data for a constant bandwidth $K(x)=I$ (red). Right: Original (blue) and generated data for a variable bandwidth $K(x)=\rho(x)I$ with $\rho(x)=\pi(x)^{\beta}$ with $\beta=-1/5$ (magenta).
Figure 3. Effect of a variable bandwidth $K(x)=\rho(x)I$ on the angular and radial distributions (left and right, respectively). Shown are the original data (blue), generated data for a constant bandwidth $K(x)=I$ (red) and for a variable bandwidth $K(x)=\rho(x)I$ with $\rho(x)=\pi(x)^{\beta}$ with $\beta=-1/5$. The data were generated using a constant covariance noise model in (21) and $\epsilon=0.009$.
Figure 4. Effect of a variable time step $\Delta\tau$ in the Langevin sampler (20) with constant diffusion $K=1$. Results are shown for the original data, and for $\Delta\tau=\epsilon$ and $\Delta\tau=\epsilon/4$. Throughout a constant bandwidth is used. Left: Empirical histogram of the angular variable $\theta$. Middle: Empirical histograms of the radial variable $r$. Right: Original (blue) and generated data in the $(x_{1},x_{2})$-plane with $\Delta\tau=1$ (red) and with $\Delta\tau=\epsilon/4$ (green).
@@ 6.2. Multi-dimensional manifolds
In this numerical example, we show our proposed method on hyper semi-spheres of dimension $d=\{3,4,9\}$, using both a fixed bandwidth kernel and a variable bandwidth kernel. Data are generated by firstly sampling $z^{(i)}=(z_{1}^{(i)},\cdots,z_{d}^{(i)})$ from a $d$-dimensional standard normal distribution, and then setting $y^{(i)}=(z^{(i)}_{1},\cdots,\alpha z^{(i)}_{d})$, with $\alpha=5$ to promote non-uniformity. Finally, the samples $x^{(i)}$ are obtained by normalizing $y^{(i)}$ to achieve the unit length, i.e., $y^{(i)}/\left\|y^{(i)}\right\|$, and perturbing $y^{(i)}/\left\|y^{(i)}\right\|$ in the radial direction with $U(0,0.01)$ noise. An instance of the target samples of three dimensions can be seen in Figure 5. Given a bandwidth $\epsilon$ and a bandwidth function $\rho(x)$, we implement the proposed scheme (21). For the fixed bandwidth kernel we set $\rho(x)=1$. For the variable bandwidth kernel (24), we set $\rho(x)=\pi(x)^{\beta}$, with $\beta<0$, and $\pi(x)$ is approximated using a kernel density estimator. We use $M=1,000$ training samples to learn the transition kernel and run a Langevin sampler to generate $50,000$ samples, with the initial data point being $(1,0,\cdots,0)\in\mathbb{R}^{d}$. To obtain a better mixing of the Langevin sampler, we take one every $20$ samples of the last $20,000$ generated samples of the chain, resulting in a total of $1,000$ samples. To evaluate the quality of the generated samples, we compute the regularized optimal transport (OT) distance between the generated samples $\{x_{\mathrm{gen}}^{(i)}\}_{i=1}^{M_{\mathrm{g}}}$ and the original reference samples $\{x_{\mathrm{ref}}^{(j)}\}_{j=1}^{M_{\mathrm{r}}}$. The regularized OT distance with entropy penalty $1/\lambda$ is defined as
@@
is the information entropy. The entries of $C\in\mathbb{R}^{M_{g}\times M_{r}}$ are set to be the pairwise Euclidean distances between $\{x_{\mathrm{gen}}^{(i)}\}_{i=1}^{M_{\mathrm{g}}}$ and $\{x_{\mathrm{ref}}^{(j)}\}_{j=1}^{M_{\mathrm{r}}}$, and each sample is assigned equal weight marginally. We compute the OT distance using the Sinkhorn–Knopp algorithm [9, 27]. The number of reference samples is chosen to be $M_{\mathrm{r}}=50,000$, and the entropic regularization penalty is set to be $1001/\lambda=$ 100. We consider the OT distance as a diagnostic to quantify the statistical accuracy of the sampling scheme.
@@
We then optimize over the parameters $\epsilon$ for a fixed $\lambda$ using grid search. To be more precise, for the Langevin sampler with a fixed bandwidth kernel, we vary $\epsilon$, and compute the OT distance of the generated samples. The best performed $\epsilon$ is chosen to be the one that corresponds to the smallest OT distance, and we call it $\epsilon^{*}$. For the variable bandwidth kernel, we fix $\epsilon=\epsilon^{*}$. In order to disentangle the effect of varying $\epsilon$ and varying $\beta$ we use the normalized variable bandwidth as described in (24). We set $\beta=-0.01\times 2^{n}$ with $n=\{0,\cdots,8\}$ for $d=\{3,4,9\}$. The optimal bandwidth $\epsilon^{*}$ is reported in Table 1 and the comparisons between the fixed bandwidth kernel and the variable bandwidth kernel are presented in Figure 6. We observe that by keeping $\epsilon$ fixed, the OT distance becomes smaller for a wide range of $\beta$.
@@
We then examine the quality of generated samples at the optimal $\epsilon$ and $\beta$ along the last coordinate (the nonuniform direction). Similar to the previous study, we compute the one dimensional OT distance of the marginal distribution (see Figure 6 (right)) and show the histograms and the cumulative density function (CDF) of the samples generated using the fixed bandwidth kernel and using the variable bandwidth kernel in Figure 7. The benefit of using the variable bandwidth kernel becomes more prominent when focusing on the marginal samples. In the case where we sample a 4-dimensional hyper-semisphere with non-uniformity along the last coordinate, we see in Figure 7 that the empirical CDF of the samples generated with the variable bandwidth kernel closely aligns with the reference (constructed using samples from the target distribution) for the most part. In contrast, the samples generated using the fixed variable bandwidth kernel noticeably diverge from the reference. This aligns with what we observed in Figure 2 and Figure 3 — the data generated using the variable bandwidth kernel better resemble the original data. In the cases where the samples are drawn from a 9-dimensional hyper-semisphere, while both methods struggle to generate samples that mirror those from the target distribution, primarily due to the inherent limitations of kernel methods in high-dimensional scenarios, employing a variable bandwidth kernel produces samples that exhibit a closer resemblance to those from the target distribution.
@@
Figure 5. Nonuniform samples $x^{(i)}$ on a $2$-dimensional semisphere. The non-uniformity is along the last coordinate $x_{3}$.
@@
Figure 6. The straight dashed lines denote the OT distances of the samples using a fixed bandwidth kernel, and the solid lines denote the OT distances of the samples using a variable bandwidth kernel. Left: comparison between the OT distance of samples generated using a fixed bandwidth kernel and using a variable bandwidth kernel. Right: comparison between the one-dimensional marginal OT distance of samples generated using a fixed bandwidth kernel and using a variable bandwidth kernel, along the last coordinate (non-uniform direction). Here the Langevin sampler is initialized at $(1,0,\cdots,0)$ for all cases.
@@
6.3. Stochastic subgrid-scale parametrization
The conditional sampling algorithm described in Section 5 can be used to perform stochastic subgrid-scale parametrization, a central problem encountered in, for example, the climate sciences. The problem of subgrid-scale parametrization, or more generally of model closure, is the following: given a potentially stiff dynamical system
@@ 6.3. Stochastic subgrid-scale parametrization
where $\varepsilon<1$ denotes the time scale separation between the slow resolved variables of interest $z\in{\mathbb{R}}^{d_{s}}$ and the unresolved fast degrees $y\in{\mathbb{R}}^{d_{f}}$. Note the notational difference between the time scale separation parameter $\varepsilon$ and the bandwidth parameter $\epsilon$ used to define the diffusion map. For $\varepsilon\ll 1$ the system is stiff and to ensure numerical stability a small time step $\Delta t<\varepsilon$ is needed. This, together with the potential high-dimensionality of the fast subspace $d_{f}\gg d_{s}$, constitutes a computational barrier for simulating the dynamics on the slow time scale of interest. Hence one is interested in obtaining an effective evolution equation for the slow resolved variables $z$ only which captures the essential effect of the unresolved variables $y$. Hence, we seek to determine the effective reduced dynamics
$\dot{z}=F_{z}(z)+\psi(z).$
Here $F_{z}(z)$ denotes a deterministic drift which we assume to be known a priori, possibly based on physical reasoning. The term $\psi(z)$ denotes the unknown closure term to be learned which may be deterministic or stochastic, and which parametrizes the unknown unresolved fast processes. Deterministic machine learning methods have previously been used to learn $\psi(z)$ as the average effect of the unresolved variables, i.e. the average of $g(z,y;\varepsilon)$ over the (conditional) invariant measure of the fast process [19, 20]. Deterministic maps, however, are not able to capture the resolved dynamics with sufficient statistical accuracy, and it is by now well established that the effective equation is often of a stochastic nature [37, 18, 25, 17]. In the case of infinite time scale separation there are explicit expressions for the effective drift and diffusion term of the effective slow dynamics. However, these terms include integrals over auto-correlation functions and are notoriously hard to estimate. Instead we propose to learn the closure term $\psi(z)$ and generate realisations $\psi(z)$ on the fly employing the conditional sampling algorithm described in Section 5. We consider the situation in which scientists have a good understanding of the resolved dynamics and know the slow vector field $F_{z}(z)$. Given data of the resolved variables $\{z_{n}\}_{n=1}^{N}$ sampled at equidistant times $\Delta t$, the closure term $\psi(z)$ capturing the effect of the unresolved dynamics (29b) can then be estimated as
$\psi(z_{n})=z_{n+1}-z_{n}-F_{z}(z_{n})\,\Delta t,$
@@
(31a) $z^{\ast}$ $=z_{n}$
(31b) $x_{1,j}$ $=z^{\ast}$
(31c) $x_{j+1/2}$ $=x_{j}+\sqrt{2\epsilon}\,\Xi_{j}$
(31d) $x_{j+1}$ $=\mathcal{Z}p_{\epsilon}(x_{j+1/2}),$
@@
We look at the case of effective additive noise with $h(z)=1$ which does not require conditioning on the slow variable $z$, as well as at the case of multiplicative noise with $h(z)=z$. We used MATLAB’s built-in ode45 routine [35] to generate the time series with a time-scale separation parameter of $\varepsilon=0.01$. The time series is subsequently sub-sampled with $\Delta t=0.1$. Figure 8 shows a comparison between the full multi-scale system (32) - (33) with additive noise $h(z)=1$ and the stochastic surrogate model (30). The slow $z$-dynamics exhibits stochastic bimodal dynamics. Figure 9 shows a comparison for the multiplicative case $h(z)=z$ which yields unimodal slow dynamics. It is clearly seen that the surrogate model (30) obtained by the generative conditional sampler generates statistically reliable dynamics.
Figure 8. Results for the stochastic subgrid-scale parametrization for the multi-scale system (32) - (33) with $\varepsilon=0.01$ and with additive noise $h(z)=1$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (31) trained with $M=120,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
Figure 9. Results for the stochastic subgrid-scale parametrization for the multi-scale system (32) - (33) with $\varepsilon=0.01$ and with multiplicative noise $h(z)=z$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (31), trained with $M=20,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
7. Summary and outlook
We introduced a Schrödinger bridge based Langevin scheme for generative modeling. Our method combines sample-based Markov chain approximations with discrete-time Langevin-type sampling, yielding a nonparametric generative model that is unconditionally stable and geometrically ergodic. We showed numerically that employing a variable bandwidth kernel, in contrast to a fixed bandwidth kernel, results in generated samples with enhanced accuracy. The performance of the conditional generative model was showcased through its application to a stochastic subgrid-scale parametrization problem.
We considered here a uniform measure on the samples. If there is a change of measure due to, for example, observed data with a given likelihood, then the measure and the couplings can be appropriately modified to take into account the non-uniform measure.
Future research will delve into the theoretical foundations of the proposed scheme, including its convergence rate and scalability. Finally, the proposed methodology can be extended to score generative modeling by replacing all required score functions by Schrödinger bridge-based approximations and to interacting particle approximations of the Fokker–Planck equation, where the score function represents now diffusion [33].